OpenStudy (anonymous):

what is the slope of 10x=9y+36?

OpenStudy (anonymous):

write equation in slope standar from y=(10x-36)/9 y=(10/9)x -4 slope=10/9

OpenStudy (anonymous):

so what is the y intercept?

OpenStudy (radar):

Review your previous problem you left early. Nancy Lam showed you the slope form of the equation. Did you not see the "-4" in her solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=-4

OpenStudy (radar):

That is the intercept, the coefficient of x is the slope.

OpenStudy (anonymous):

thanks

OpenStudy (radar):

Yes y=-4 when x=0